# Optimizing Sequencing Coverage Depth in DNA Storage: Insights From DNA Storage Data

### v1

@@ Abstract
DNA data storage is now being considered as a new archival storage method for its durability and high information density, but still facing some challenges like high costs and low throughput. By reducing sequencing sample size for decoding digital data, minimizing DNA coverage depth helps lower both costs and system latency. Previous studies have mainly focused on minimizing coverage depth in uniform distribution channels under theoretical assumptions. In contrast, our work uses real DNA storage experimental data to extend this problem to log-normal distribution channels, a conclusion derived from our PCR and sequencing data analysis. In this framework, we investigate both noiseless and noisy channels. We first demonstrate a detailed negative correlation between linear coding redundancy and the expected minimum sequencing coverage depth. Moreover, we observe that the probability of successfully decoding all data in a single sequencing run increases and then decreases as coding redundancy rises, when the sample size is optimized for complete decoding. Then we extend the lower bounds of DNA coverage depth from uniform to log-normal noisy channels. The findings of this study provide valuable insights for the efficient execution of DNA storage experiments.
@@
DNA data storage, sequencing coverage depth, log-normal distribution channel.
@@ I Introduction
With the rapid growth of global data and advancements in information technology, traditional storage media such as HDDs and SSDs may no longer meet future data storage demands [1, 2]. DNA storage, with its high density and durability, has emerged as a promising solution to address this challenge. However, current DNA storage technologies face significant obstacles, including system latency due to low throughput and high costs. Minimizing sequencing coverage depth in DNA storage systems could provide valuable insights for reducing latency and associated costs.
Several studies have addressed the issue of sequencing coverage depth in non-random access scenarios, focusing on minimizing sequencing coverage depth in channels follow uniform distribution by adjusting outer error-correcting codes. These approaches are often based on classical problems such as the coupon collector’s problem, the urn problem, and others. [3] first introduced the concept of sequencing coverage depth, investigating the minimum coverage depth required for successful message decoding under non-random access settings. They also explored the singleton coverage depth problem and proposed a new coding scheme to reduce the sample size needed to retrieve a single strand. [4] extended the work of [3] to more practical application scenarios, particularly focusing on retrieving $k$ (where $k=1$ or $2$) out of $n$ files. This study also examined the structural properties of various coding schemes and their impact on random access expectations and probability distributions. [5] provided a novel perspective on random access efficiency by exploring coding matrix design and performance optimization. Their work introduced the concept of recovery balanced codes and corresponding conditions, offering theoretical guidance for designing more efficient DNA storage systems. [6] proposed the first model to analyze sequencing coverage depth in combinatorial coding, calculating the probability of error-free reconstruction and investigating theoretical bounds to evaluate decoding probabilities. These studies collectively address the problem of minimizing sequencing coverage depth in uniformly distributed channels, offering various theoretical models, coding schemes, and analytical methods. Their contributions provide a foundation for optimizing sequencing efficiency and cost-effectiveness in DNA storage systems. However, the processes of DNA synthesis and amplification exhibit a degree of randomness, which results in a non-uniform channel probability distribution. Therefore, in this study, we investigate the problem of minimizing sequencing coverage depth in a real-world channel based on PCR and sequencing data, under the non-random access setting.
@@ II-A Problem Statement
For alphabet $\varSigma=\{A,C,G,T\}$, each DNA strand or sequence of length $l$ can be represented as a vector of length $l$ over the alphabet $\varSigma$. The set of all such vectors of length $l$ over $\varSigma$ is denoted by $\varSigma^{l}$. Let $\varSigma^{*}\triangleq\bigcup_{l=0}^{\infty}\varSigma^{l}$.
@@
Write and encode. Assume that the digital data to be stored is first segmented and translated into $m$ information strands of length $l$, denoted as $X=(x_{1},x_{2},\dots,x_{m})\in(\varSigma^{l})^{m}$. After encoding with error-correcting code $\mathcal{C}$, the $m$ information strands are transformed into $n$ designed strands of length $l$, represented as $D=(d_{1},d_{2},\dots,d_{n})\in(\varSigma^{l})^{n}$. Next, each designed strand in $D$ is synthesized into $c$ copies. Due to the inherent randomness in the synthesis process, the actual copy numbers of each designed strand $c_{i}$ is not the same [12], where $i\in[1,n]$, and each synthesis strand is a noisy version of its corresponding designed strand. Let $M=\sum_{i=1}^{n}c_{i}$, then, all of $M$ synthetic strands will be stored unordered in a container until reading them out.
@@
Read and decode. In this step, we will extract a specific volume or weight of DNA dry powder in the container and performing sequencing using next-generation sequencer (NGS). Let the multiple set $S_{K}=\{\{s_{1},s_{2},\dots,s_{K}\}\}$ represents the $K$ reads obtained from sequencing, where each $s_{i}\in\varSigma^{*}$ is a noisy copy of a particular designed strand $d_{j}$. During the sequencing process, due to differences in sequencing technologies, the reads in $S_{K}$ can either be obtained sequentially, as in Nanopore sequencing, or all at once, as in Illumina sequencing. However, the research conclusions drawn from both approaches do not show significant differences [7]. In this study, a sequential reading method is employed. The number of noisy copies of the $i$-th designed strand $d_{i}$ in $S_{K}$ depends on the channel probability distribution $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, where $p_{i}^{(t)}$ represents the probability of obtaining a copy of $d_{i}$ in each round, i.e., the proportion of $d_{i}$ in the population after certain cycles of PCR amplification. During the decoding process, we first apply a clustering algorithm to partition the reads in $S_{K}$ into clusters based on their corresponding designed strands. Specifically, the $i$-th cluster contains all noisy copies of the $i$-th designed strand $d_{i}$ in $S_{K}$, denoted as $s_{J}=\{s_{j}|s_{j}\ \text{is a noisy copy of}\ d_{j}\}$. Based on the assumptions and analysis outlined above, the probability of successfully retrieving both the information strand set $X$ and the designed strand set $D$ primarily depends on the error-correcting code $\mathcal{C}$ and the retrieval algorithm [7].
@@
    Explanation of error-correcting code $\mathcal{C}$. The construction of the error-correcting code determines how many designed strands need to be retrieved to decode all the digital data. In our work, for given $m$ information strands, we assume the use of a $[n,m]$ MDS code $\mathcal{C}$ with redundancy $R=\frac{n}{m}$, meaning that during the decoding process, retrieving only $m$ out of $n$ designed strands is sufficient to decode the entire dataset.
@@
    Explanation of retrieval algorithm. The retrieval algorithm determines how many noisy copies of a designed strand need to be retrieved in order to successfully decode that strand. In our work, we model the retrieval algorithm using a positive integer $a\geq 1$, assuming that at least $a$ noisy copies of $d_{i}$ need be retrieved for $d_{i}$ to be successfully decoded.
To this end, we model the sample size as a function of the channel probability distribution $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, the [n,m] MDS code $\mathcal{C}$, and the retrieval algorithm $a$, where $p_{i}$ represents the probability that a single read is from the $i$-th designed strand after $t$ cycles of PCR amplification, and $a$ is a positive integer. We define a new notation $K\triangleq I_{a}^{\boldsymbol{p}_{t}}(n,m)$, which means the sample size $K$ to decode the full digital data under the condition that after $t$ cycles of PCR, the retrieval algorithm needs to retrieve at least $a$ noisy copies of each of $m$ out of $n$ designed strands. For a uniformly distributed channel, we will omit $\boldsymbol{p}_{t}$, and denote it as $I_{a}(n,m)$.
The main problems we investigate in this paper are defined below.
@@
Problem 2. (Coverage depth problem in the real channel.) Given values of $n\geq m\geq 1$, $a\geq 1$, we focus on the following questions:
@@
    The expectation value $\mathbb{E}[I_{a}(n,m)]$.
@@
    Given $t\geq 1$, the expectation value $\mathbb{E}[I_{1}^{\boldsymbol{p}_{t}}(n,m)]$.
@@
    Given $t\geq 1$, the lower bounds of $I_{a}^{\boldsymbol{p}_{t}}(n,m)$.
@@ II-B Previous Work
$\mathbb{E}[I_{1}(n,m=n)]=n\log n+\gamma n+\mathcal{O}(1),$
@@
$\mathbb{E}[I_{a}^{\boldsymbol{p}}(n,m)]=\sum_{q=0}^{m-1}\int_{0}^{\infty}[v^{q}]\prod_{i=1}^{n}(e_{t-1}(p_{i}r)+v(e^{p_{i}r}-e_{t-1}(p_{i}r)))e^{-r}dr,$ (1)
where $\boldsymbol{p}$ represents any general discrete distribution, $[v^{q}]Q(v)$ represents the coefficients of the $q$-th terms of the polynomial $Q(v)$, and $e_{t}(x)=\sum_{i=1}^{t}\frac{x^{i}}{i!}$.
@@
$\log\left(\frac{1}{1-R}\right)+f_{c}(n,R)\leq\mathbb{E}\left[\frac{I_{a}(n,m)}{n}\right]\leq\left(\log\left(\frac{1}{1-R}\right)+a\log\log n+2\log(a+1)\right)\cdot(1+2\epsilon),$
@@
Obviously, (1) is not a closed form expression and is not easy to calculate. Thus, we will build upon the MDS coverage depth problem in [3] that addresses the minimum coverage depth in uniform distribution channels under non-random access setting to derive the expectation value of $I_{a}^{\boldsymbol{p}_{t}}(n,m)$ in a noiseless channel, as well as its lower bounds in a noisy channel in Section IV.
@@ III Simulation of DNA Storage Channel Probability Distribution
According to the definition of the channel probability distribution in Section II, it is closely related to the proportion of each strand in the population after PCR amplification. In this section, by analyzing PCR and sequencing data, we find that the real channel probability distribution follows a log-normal distribution. Based on this, we propose a theoretical model for the simulation of the real channel probability distribution, which serves as the foundation for the subsequent research on minimizing sequencing coverage depth.
@@ III-A Analysis of Channel Probability Distribution Based on PCR and Sequencing Data
We first analyze the real channel probability distribution based on the PCR and sequencing data. We have an oligo pool of 11,520 distinct oligos(i.e., $n=11520$), and PCR amplification was performed using these oligos as template strands. After 10 cycles of amplification, sequencing 4,970,786 reads to generate Dataset PCR10; after 30 cycles, sequencing 11,001,029 reads to generate Dataset PCR30; and after 60 cycles, sequencing 11,180,177 reads to generate Dataset PCR60. The relationship between different PCR cycles and the copy numbers of different oligos at that cycle is shown in Fig. 4(a), 4(b) and 4(c).
@@
Our primary interest lies in the relationship between the number of PCR cycles and the proportion of each strand in the population. Given that the sequencing data exhibit a clear skewed distribution, we assume that the PCR data follow a log-normal distribution [13]. Then we performed a logarithmic transformation on the normalized data before fitting a normal distribution, and Fig. 4(d), 4(e) and 4(f) show the results of a good fit on all three datasets, which means after PCR amplification, the channel actually follows a log-normal distribution.
@@ III-B Real Channel Probability Distribution Model
In the previous subsection, we analyzed PCR experimental data to obtain the real channel distributions after 10, 30, and 60 cycles of PCR amplification. However, in practical experiments, additional sequencing of PCR data to obtain the parameter information of the probability distribution is typically not performed. This necessitates the simulation of the channel probability distribution under the condition that only the synthetic amount $c_{i}$ and amplification efficiency $r_{i}$ of each designed strand are known.
Therefore, for given $c_{i}$ and $r_{i}$, where $i\in[1,n]$, we model the channel probability distribution as a function of PCR cycles. First, it is evident that after $t$ cycles of PCR amplification, the expected copy number of the $i$-th strand is
$\mathbb{E}\left[\nu_{i}^{(t)}\right]=c_{i}(1+r_{i})^{t}.$
@@
then the expectation and variance of $p_{i}^{(t)}$ can be calculated, denoted by $\mathbb{E}\left[p_{i}^{(t)}\right]$ and $Var\left[p_{i}^{(t)}\right]$.
@@
$\mu^{(t)}=\log{\left(\mathbb{E}\left[p_{i}^{(t)}\right]\right)}-\frac{1}{2}\log{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)},$ (4)
@@
$\sigma^{(t)}=\sqrt{\ln{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)}}.$ (5)
@@
$p_{i}^{(t)}\sim\mathcal{LN}\left(\log{\left(\mathbb{E}\left[p_{i}^{(t)}\right]\right)}-\frac{1}{2}\log{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)},\ln{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)}\right).$
@@
IV Sequencing Coverage Depth Problem in the Log-normal Distribution Channel
Under the condition of non-random access, the main goal of our study is to find the minimum sequencing sample size $K$ that ensures the successful decoding of all digital data in the real channel that following the log-normal distribution after PCR. In this section, we first prove the expected coverage depth in log-normally distributed noiseless channels, and propose the problem of decoding all data successfully in a single sequencing experiment using the expected coverage depth as the sample size. Then we prove the theoretical lower bounds of expected coverage depth in the real noisy channels. For given $m$ information strands, we denote $\alpha\triangleq\frac{K}{m}$ as the sequencing coverage depth.
IV-A Sequencing Coverage Depth in the Noiseless Channel
In practical experiments, in addition to focusing on the expected sequencing coverage depth required to decode all the digital information (i.e., $\mathbb{E}[\alpha]$), we are also concerned with the probability of successfully decoding all data in a single sequencing run (i.e., $Var[\alpha]$). Let $N$ represents the number of different designed strands when sequencing fixed $K$ reads.

### v2

@@ Abstract
DNA storage is now being considered as a new archival storage method for its durability and high information density, but still facing some challenges like high costs and low throughput. By reducing sequencing sample size for decoding digital data, minimizing DNA coverage depth helps lower both costs and system latency. Previous studies have mainly focused on minimizing coverage depth in uniform distribution channels under theoretical assumptions. In contrast, our work uses real DNA storage experimental data to extend this problem to log-normal distribution channels, a conclusion derived from our PCR and sequencing data analysis. In this framework, we investigate both noiseless and noisy channels. We first demonstrate a detailed positive correlation between MDS code rate and the expected minimum sequencing coverage depth. Moreover, we observe that the probability of successfully decoding all information in a single sequencing run decreases and then increases as code rate rises, when the sample size is optimized for complete decoding. Then we extend the lower bounds of the DNA coverage depth from uniform to log-normal noisy channels. The findings of this study provide valuable insights for the efficient execution of DNA storage experiments.
@@
DNA storage, sequencing coverage depth, log-normal distribution channel.
@@ I Introduction
With the rapid growth of global data and advancements in information technology, traditional storage media such as HDDs and SSDs may no longer meet future data storage demands [1, 2]. DNA storage, with its high density and durability, has emerged as a promising solution to address this challenge. However, current DNA storage technologies face significant obstacles, including low throughput and high costs. Reducing the sequencing sample size required to ensure a high probability of decoding all information is the main goal of coverage depth problem, which could provide valuable insights for reducing latency and associated costs.
Several studies have addressed this issue in DNA storage channels follow uniform distribution by adjusting outer error-correcting codes. These approaches are often based on classical problems such as the coupon collector’s problem[3], the urn problem[4], and others. [5] first introduced the concept of sequencing coverage depth, investigating the minimum coverage depth required for successful message decoding under non-random access settings. They also explored the singleton coverage depth problem and proposed a new coding scheme to reduce the sample size needed to retrieve a single strand. [6] extended the work of [5] to more practical application scenarios, particularly focusing on retrieving $k$ (where $k=1$ or $2$) out of $n$ files. This study also examined the structural properties of various coding schemes and their impact on random access expectations and probability distributions. [7] provided a novel perspective on random access efficiency by exploring coding matrix design and performance optimization. Their work introduced the concept of recovery balanced codes and corresponding conditions, offering theoretical guidance for designing more efficient DNA storage systems. [8] proposed the first model to analyze sequencing coverage depth in combinatorial coding, calculating the probability of error-free reconstruction and investigating theoretical bounds to evaluate decoding probabilities. [9] introduced a new geometric structure, namely the balanced quasi-arc, to study the random access problem. By analyzing this geometric structure, the study aimed to identify encoding schemes that could effectively reduce the expected number of samples required for random access.
These studies collectively address the problem of minimizing sequencing coverage depth in uniformly distributed channels, offering various theoretical models, coding schemes, and analytical methods. Their contributions provide a foundation for optimizing sequencing efficiency and cost-effectiveness in DNA storage systems. However, the processes of DNA synthesis and amplification exhibit a degree of randomness, which results in a non-uniform channel probability distribution. Therefore, in this study, we investigate the problem of minimizing sequencing coverage depth in a real-world channel based on PCR (Polymerase Chain Reaction) and sequencing data, under the non-random access setting.
@@ II-A Problem Statement
For alphabet $\varSigma=\{A,C,G,T\}$, each DNA strand of length $l$ can be represented as a vector of length $l$ over the alphabet $\varSigma$. The set of all such vectors of length $l$ over $\varSigma$ is denoted by $\varSigma^{l}$. Let $\varSigma^{*}\triangleq\bigcup_{l=0}^{\infty}\varSigma^{l}$, and $[1,n]=1,2,\dots,n$.
@@
Write and encode. Assume that the digital data to be stored is first segmented and translated into $m$ information strands of length $l$, denoted as $X=(x_{1},x_{2},\dots,x_{m})\in(\varSigma^{l})^{m}$. After encoding with an $[n,m]$ MDS code[10], the $m$ information strands are transformed into $n$ designed strands of length $l$, represented as $D=(d_{1},d_{2},\dots,d_{n})\in(\varSigma^{l})^{n}$. Next, each designed strand in $D$ is synthesized into $c$ copies. Due to the inherent randomness in the synthesis process, the actual copy numbers of each designed strand $c_{i}$ is not the same [11], where $i\in[1,n]$, and each synthesis strand is a noisy version of its corresponding designed strand. Then all synthetic strands will be stored unordered in a container until reading them out.
@@
Read and decode. In this step, we will extract a specific volume or weight of DNA dry powder in the container and performing sequencing using next-generation sequencer (NGS). Let the multiple set $S_{K}=\{\{s_{1},s_{2},\dots,s_{K}\}\}$ represents the $K$ reads obtained from sequencing, where each $s_{i}\in\varSigma^{*}$ is a noisy copy of a particular designed strand $d_{j}$. The number of noisy copies of the $i$-th designed strand $d_{i}$ in $S_{K}$ depends on the probability distribution $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, where $p_{i}^{(t)}$ represents the probability of sampling a read of $d_{i}$ in each round, i.e., the proportion of $d_{i}$ in the population after $t$ cycles of PCR amplification. During the decoding process, we first apply a clustering algorithm to partition the reads in $S_{K}$ into clusters based on their corresponding designed strands. Specifically, the $i$-th cluster contains all noisy copies of the $i$-th designed strand $d_{i}$ in $S_{K}$, denoted as $s_{J}=\{s_{j}|s_{j}\ \text{is a noisy copy of}\ d_{j}\}$. Then we use the retrieved designed strands to recover the rest information strands.
Remark. During the sequencing process, due to differences in sequencing technologies, the reads in $S_{K}$ can either be obtained sequentially, as in Nanopore sequencing, or all at once, as in Illumina sequencing. However, the research conclusions drawn from both approaches do not show significant differences [5]. In this study, a sequential reading method is employed.
We formulate this coverage depth problem under non-random access setting as a variant of the classical coupon collector’s problem or the urn model. Specifically, we consider the scenario in which one identical ball is thrown in each round, and the probability of a ball falling into each urn varies. After throwing $K$ rounds, we are interested in the number of urns—among $n$ indistinguishable urns except for their labels—that contain at least $a$ balls. To this end, we model the sample size as a function of the channel probability distribution, the MDS code, and the retrieval algorithm. We proceed to give a detailed explanation of these three variables.
@@
    Explanation of the channel probability distribution. The channel probability distribution is analogous to the probability distribution of a ball falling into each urn. Unlike [5], which assumes a uniform distribution of the number of each designed strands, i.e., the probability of a ball falling into each urn is equal, we consider that the randomness in synthesis and PCR amplification causes varying copy numbers of each designed strand, leading to unequal proportions in the population. We assume that the number of noisy copies of each designed strand observed in the sequencing reads follows a probability distribution $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, where $p_{i}^{(t)}$ denotes the probability of sampling a read corresponding to the $i$-th strand after $t$ cycles of PCR amplification. Notice that for simplicity, we consider the distribution $\boldsymbol{p}_{t}$ solely as a function of the DNA storage channel, without accounting for potential influences from strand design [5]. Accordingly, we refer to $\boldsymbol{p}_{t}$ as the channel probability distribution in this paper.
@@
    Explanation of the MDS code. The construction of the error-correcting code determines how many designed strands need to be retrieved to decode all digital data, i.e., how many urns are needed to contain at least $a$ balls after $K$ rounds? In our work, when an $[n,m]$ MDS code is used to encode the information strands, successful retrieval of any $m$ out of the $n$ encoded strands is sufficient to fully recover the original information. We denote the code rate of the $[n,m]$ MDS code by $R=\frac{m}{n}$.
  3. 3.
    Explanation of the retrieval algorithm. The probability of successful retrieval of each designed strand depends on the number of its noisy copies in the sequencing pool, the use of inner codes during encoding, and the channel error rate [5]. In this paper, we model the retrieval algorithm by introducing a positive integer parameter $a\geq 1$, assuming that each designed strand can be successfully retrieved if at least $a$ reads corresponding to it are available.
Based on the assumptions and analysis outlined above, the main problems we investigate in this paper are defined below.
@@
Problem 2. (MDS coverage depth problem in the real channel.) Given values of $n\geq m\geq 1$, $a\geq 1$, we focus on the following questions:
@@
    The expectation value $\mathbb{E}[K_{a}(n,m)]$.
@@
    Given $t\geq 1$, the expectation value $\mathbb{E}[K_{1}^{\boldsymbol{p}_{t}}(n,m)]$.
@@
    Given $t\geq 1$, the lower bounds of $K_{a}^{\boldsymbol{p}_{t}}(n,m)$.
@@ II-B Previous Work
$\mathbb{E}[K_{1}(n=m,m)]=m\log m+\gamma m+\mathcal{O}(1),$
@@
$\mathbb{E}[K_{a}^{\boldsymbol{p}}(n,m)]=\sum_{q=0}^{m-1}\int_{0}^{\infty}[v^{q}]\prod_{i=1}^{n}(e_{t-1}(p_{i}r)+v(e^{p_{i}r}-e_{t-1}(p_{i}r)))e^{-r}dr,$ (1)
where $\boldsymbol{p}$ represents any general discrete distribution, $[v^{q}]Q(v)$ represents the coefficients of the $q$-th terms of the polynomial $Q(v)$, and $e_{t}(x)=\sum_{i=0}^{t}\frac{x^{i}}{i!}$.
@@
$\log\left(\frac{1}{1-R}\right)+f_{c}(n,R)\leq\mathbb{E}\left[\frac{K_{a}(n,m)}{n}\right]\leq\left(\log\left(\frac{1}{1-R}\right)+a\log\log n+2\log(a+1)\right)\cdot(1+2\epsilon),$
@@
Equation (1) does not provide a closed-form expression and is not straightforward to compute. Thus, we will build upon the MDS coverage depth problem in [5] that addresses the minimum coverage depth in uniform distribution channels under non-random access setting to derive the expectation value of $K_{a}^{\boldsymbol{p}_{t}}(n,m)$ in a noiseless channel, as well as its lower bounds in a noisy channel in Section IV.
@@ III Simulation of DNA Storage Channel Probability Distribution
According to the definition of the channel probability distribution in Section II, it is closely related to the proportion of each designed strand in the population after PCR amplification. In this section, by analyzing PCR and sequencing data, we find that the real channel probability distribution follows a log-normal distribution. Based on this, we propose a theoretical model for the simulation of the real channel probability distribution, which serves as the foundation for the subsequent research on minimizing sequencing coverage depth.
@@ III-A Analysis of Channel Probability Distribution Based on PCR and Sequencing Data
We first analyze the real channel probability distribution based on the PCR and sequencing data. We synthesized 11,520 oligos, each 150 bases long, using inkjet printing technology at Twist Bioscience (i.e., $n=11520$), and PCR amplification was performed using these oligos as template strands. After 10 cycles of amplification, sequencing 4,970,786 reads to generate Dataset PCR10; after 30 cycles, sequencing 11,001,029 reads to generate Dataset PCR30; and after 60 cycles, sequencing 11,180,177 reads to generate Dataset PCR60. The relationship between different PCR cycles and the copy numbers of different oligos at that cycle is shown in Fig. 2(a), 2(b) and 2(c).
@@
Our primary interest lies in the relationship between the number of PCR cycles and the proportion of each designed strand in the population. Given that the sequencing data exhibit a clear skewed distribution, we assume that the PCR data follow a log-normal distribution [16]. Then we performed a logarithmic transformation on the normalized data before fitting a normal distribution, and Fig. 2(d), 2(e) and 2(f) show the results of a good fit on all three datasets, which means after PCR amplification, the channel actually follows a log-normal distribution.
@@ III-B Real Channel Probability Distribution Model
In the previous subsection, we analyzed PCR experimental data to obtain the real channel distributions after 10, 30, and 60 cycles of PCR amplification. However, in practical experiments, additional sequencing of PCR data to obtain the parameter information of the probability distribution is typically not performed. This necessitates the simulation of the channel probability distribution under the condition that only the synthetic amount $c_{i}$ and amplification efficiency $r_{i}$ of each designed strand are known, where PCR amplification efficiency is the rate at which the target DNA fragment is amplified during each cycle in the exponential phase. Theoretically, perfect replication would result in 100% efficiency, doubling the DNA amount per cycle. However, in practice, efficiency is typically below 100% and ranges from 80% to 110% due to various factors.
Therefore, for given $c_{i}$ and $r_{i}$, where $i\in[1,n]$, we model the channel probability distribution as a function of PCR cycles. We first provide the calculation process for the expected copy number of the $i$-th strand after $t$ cycles of PCR amplification, i.e., $\nu_{i}^{(t)}$.
According to the definition of amplification efficiency, let $\mathbb{E}\left[\nu_{i}^{(0)}\right]=c_{i}$, for $t=1$:
$\mathbb{E}\left[\nu_{i}^{(1)}\right]=\mathbb{E}\left[\nu_{i}^{(0)}\right]\cdot(1+r_{i})=c_{i}(1+r_{i}).$
For $t=2$:
$\mathbb{E}\left[\nu_{i}^{(2)}\right]=\mathbb{E}\left[\nu_{i}^{(1)}\right]\cdot(1+r_{i})=c_{i}(1+r_{i})^{2}.$
And so on, for $t=q$:
$\mathbb{E}\left[\nu_{i}^{(q)}\right]=\mathbb{E}\left[\nu_{i}^{(q-1)}\right]\cdot(1+r_{i})=c_{i}(1+r_{i})^{q}\quad\text{for all positive integers q}.$
@@
then the expectation and variance of $p_{i}^{(t)}$ can be calculated, denoted by $\mathbb{E}\left[p_{i}^{(t)}\right]$ and $\mathrm{Var}\left[p_{i}^{(t)}\right]$.
@@
$\mu^{(t)}=\log{\left(\mathbb{E}\left[p_{i}^{(t)}\right]\right)}-\frac{1}{2}\log{\left(1+\frac{\mathrm{Var}\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)},$ (4)
@@
$\sigma^{(t)}=\sqrt{\ln{\left(1+\frac{\mathrm{Var}\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)}}.$ (5)
@@
$p_{i}^{(t)}\sim\mathcal{LN}\left(\log{\left(\mathbb{E}\left[p_{i}^{(t)}\right]\right)}-\frac{1}{2}\log{\left(1+\frac{\mathrm{Var}\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)},\ln{\left(1+\frac{\mathrm{Var}\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)}\right).$
@@
IV MDS Coverage Depth Problem in the Log-normal Distribution Channel
Under the condition of non-random access, the main goal of our study is to find the minimum sequencing sample size $K$ that ensures the successful decoding of all digital data in the real channel that follows the log-normal distribution after PCR. In this section, we first prove the expected coverage depth in log-normally distributed noiseless channels, and propose the problem of decoding all data successfully in a single sequencing experiment under the expected sample size. Then we prove the theoretical lower bounds of expected sample size in the real noisy channels. For given $n$ designed strands, we denote $\alpha\triangleq\frac{K}{n}$ as the sequencing coverage depth.
IV-A MDS Coverage Depth Problem in the Noiseless Channel
In practical experiments, in addition to focusing on the expected sequencing coverage depth required to decode all digital information (i.e., $\mathbb{E}[\alpha]$), we are also concerned with the probability of successfully decoding all data in a single sequencing run (i.e., $\mathrm{Var}[\alpha]$). Let $N$ represent the number of different designed strands when sequencing fixed $K$ reads.